OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took one-third the number of tickets that her cousin took, which was half the number Heather took. If her brother took 5 tickets, how many tickets did Heather take? fifteen thirty forty forty-five

OpenStudy (anonymous):

@bahrom7893 can you help me

OpenStudy (bahrom7893):

Let x be the number of the total tickets bought. Let y be the number of the tickets that Heather took.

OpenStudy (anonymous):

@bahrom7893 ok

OpenStudy (bahrom7893):

This means that Heather's cousin, which took half the number Heather took, took: (y/2) tickets => Heather had y tickets, her brother had half of that amount, hence y/2. So far so good?

OpenStudy (dumbcow):

we got little mermaid and dopey ...what a pair

OpenStudy (anonymous):

@dumbcow lol

OpenStudy (bahrom7893):

hahahahah.. Anyway, Juliette, so far so good?

OpenStudy (anonymous):

@bahrom7893 yes

OpenStudy (bahrom7893):

So now we know that: There are x tickets in total. Heather has y tickets. Her cousin has (y/2) tickets. Her brother has (1/3) of the amount of tickets that her cousin has, so her brother has: (1/3)*(y/2).

OpenStudy (bahrom7893):

So now we have: x - total y - Heather y/2 - cousin y/6 - brother. And we also know that her brother took 5 tickets, so: y/6 = 5

OpenStudy (anonymous):

Yes

OpenStudy (bahrom7893):

6*y/6 = 6*5 = > y = 30

OpenStudy (anonymous):

@bahrom7893 thank you!

OpenStudy (bahrom7893):

And since y was the number of tickets Heather had, she has 30 tickets.